MILLIMETER WAVE SPECTRA OF CHLORINE NITRATE

\underline{REBECCA A.H. BUTLER}, Department of Physics, The Ohio State University, 174 West 18th Avenue, Columbus, OH 43210; PAUL HELMINGER, Department of Physics and Chemistry, University of South Alabama, Mobile, AL 36688; FRANK C. DE LUCIA, Department of Physics, The Ohio State University, 174 West 18th Avenue, Columbus, OH 43210.

Chlorine nitrate is a planar, slightly asymmetric molecule of atmospheric interest. It has a dense rotational spectrum due to its relatively small rotational constants and its several low-lying vibrational modes. We have recorded this spectrum with a FASSST spectrometer over a range of 128-350 GHz, and will report on our analysis of many of these lower vibrational states, including 2ν_{9}, 3ν_{9}, 4ν_{9}, 5ν_{9}, ν_{7}, 2ν_{7}, ν_{7}+ν_{9}, and ν_{6}, taking into account several perturbations between states.